OpenStudy (anonymous):

Find all values of t in [0, 2π] for which cot^2t=3

OpenStudy (anonymous):

cot^2t=3 tan^2 t = 1/3

OpenStudy (anonymous):

Hope now u can do it

OpenStudy (anonymous):

Where did you get tangent?

terenzreignz (terenzreignz):

Tangent is just the reciprocal of cotangent.

terenzreignz (terenzreignz):

So, if cot x = a (a not being equal to zero, of course0 then tan x = 1/a

OpenStudy (anonymous):

So am I supposed to solve both?

terenzreignz (terenzreignz):

no. Whichever is easier to you, solve that. They are basically the same equation, anyway.

OpenStudy (anonymous):

I got 0.524

OpenStudy (anonymous):

Can you help with another?

terenzreignz (terenzreignz):

Wait, what? How did you get that?